OpenStudy (anonymous):

A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer?

OpenStudy (anonymous):

so confusing!

OpenStudy (anonymous):

A = (pi)[(x + 10)^2 - (x + 6)^2] A = (pi)(x^2 + 20x + 100 - x^2 - 12x - 36) A = (pi)(8x + 64) = (8)(pi)(x + 8)

OpenStudy (anonymous):

You are getting the area of the circle and you are then subtracting the area of the portion drilled out.

OpenStudy (anonymous):

oooooooh!!! i see now. ty ty ty!

OpenStudy (anonymous):

You're welcome!